# Distributionally Robust Optimization with Adversarial Data Contamination

### v1

@@ 3.3 Equivalent DRO Reformulation
In this work, we focus on convex Lipschitz-continuous functions that can be written as a composition of a high-dimensional affine function and a univariate nonlinear function. In particular, if for labeled data $({\bm{x}},y)$ the loss function $\ell({\bm{w}};{\bm{x}},y)$ takes the form of either $\phi_{1}({\bm{w}}\cdot{\bm{x}}-y)$ [^6] More generally, for loss functions of the form $\phi_{1}({\bm{w}}\cdot{\bm{x}}-y)$, we can also allow for distributional shifts in the label $y$: when $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|[{\bm{x}}^{\top},y]^{\top}-[{\bm{x}}^{\prime\top},y^{\prime}]^{\top}\|_{r}$, we replace the regularization $\|{\bm{w}}\|_{s}$ by $\|[{\bm{w}}^{\top},-1]^{\top}\|_{s}$. For conciseness, in this paper we only consider distributional shifts in the covariates ${\bm{x}}$. or $\phi_{2}(y{\bm{w}}\cdot{\bm{x}})$ for univariate functions $\phi_{1}$ and $\phi_{2}$ of Lipschitz modulus $\zeta$ and the cost function is $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}+\chi(y-y^{\prime})$ for some $r\in[1,+\infty]$, where $\chi(0)=0$ and is +$\infty$+ ∞ elsewhere, then [shafieezadeh2019regularization, Theorem 4 and Remark 18] implies that (1) is equivalent to
@@
where $11/r+1/s=$ 1 . Thus, our problem reduces to solving (3) with sample access to data-contaminated distribution ${{\mathbb{P}}_{\text{train}}}$ (cf. (2)). Due to standard uniform convergence results, it suffices to robustly solve an empirical version of (3) as in problem (P).
We give a few examples of reformulating Equation 1 as Equation 3. Recall that the cost function is $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}+\chi(y-y^{\prime})$; in the following, we drop the superscript $c$, since the context is clear. The convex conjugate of these loss functions and their efficient proximal operator evaluation per Definition 2.1 can be found in [beck2017first, Sections 4.4.16 and 6.9].
@@
Our results will apply to both Example C.1 and Example 3.2. For completeness, we also provide the following example that can be reformulated similar to (3), but to which our results will not apply, due to the lack of finite-sum structure in the loss function after reformulation.
@@ 3.4 Inexact Hybrid Gradient Oracle
Finally, we argue that under standard distributional assumptions (stated as Assumption 3.4), we can assume access to a bounded-error oracle for data drawn from ${\mathbb{P}}_{0}.$
@@
The stipulation in Assumption 3.4 that covariates $\tilde{{\bm{x}}}$ (i.e., ${\bm{x}}$ excluding its first, constant component) are centered, meaning $\mathbb{E}[\tilde{{\bm{x}}}]={\bm{0}}$, is a standard simplification that can be made without loss of generality for linear prediction models incorporating an intercept term $w^{0}$. Indeed, for a fixed $\tilde{\bm{w}}\in\mathbb{R}^{d-1}$, $\tilde{\bm{w}}\cdot\tilde{\bm{x}}+w^{0}=\tilde{\bm{w}}\cdot(\tilde{\bm{x}}-\mathbb{E}[\tilde{\bm{x}}])+w^{0}+\tilde{\bm{w}}\cdot\mathbb{E}[\tilde{\bm{x}}]$. This shows that a model prediction based on potentially uncentered covariates $\tilde{\bm{x}}$ with an intercept $w^{0}$ is identical to a prediction using centered covariates $(\tilde{{\bm{x}}}-\mathbb{E}[\tilde{{\bm{x}}}])$ if we use the exact same weight $\tilde{{\bm{w}}}$ but adjust the intercept to be $(w^{0}+\tilde{{\bm{w}}}\cdot\mathbb{E}[\tilde{{\bm{x}}}])$. The mean $\mathbb{E}[\tilde{\mathbf{x}}]$ can be estimated from $\epsilon$-corrupted data using Algorithm 3. See Section C.5 for details. We verify in Section C.3 that prepending covariates by $1$ preserves the bounded covariance assumption.
Then we prove the following key technical lemma that enables the robust estimation of inexact hybrid gradients for all iterations, which is a nontrivial task because iterates depend on previous iterations.
@@
This lemma implies that we can use Lemma 3.5 to construct an oracle that satisfies Assumption 1.4 via robust mean estimation algorithms described in Section 2.
@@
On input an arbitrary $3\zeta$-uniformly bounded sequence $(\beta^{i})_{i=1}^{N}$ and a $2\epsilon$-corrupted version of an $(O(\epsilon),O(\sqrt{\epsilon}))$-stable set $\{{\bm{x}}_{i}\}_{i=1}^{N}$ with respect to variance $\sigma^{2}$, $\mathsf{RobustMeanEstimation}$ (Algorithm 3) outputs $\hat{{\bm{z}}}$ such that $\|\hat{\bm{z}}-\frac{1}{N}\sum_{i=1}^{N}\beta^{i}{\bm{x}}_{i}\|_{2}=O(\sigma\zeta\sqrt{\epsilon})$.
@@ 4 Algorithm and Analysis
In this section, we show how to solve convex Lipschitz data-contaminated $W_{1}$-DRO problems using an efficient primal-dual algorithm. Here we crucially rely on the problem reformulation described in the previous section to reduce the problem to minimizing functions of the form $\frac{1}{N}\sum_{i=1}^{N}l_{i}({\bm{x}}_{i}\cdot\hat{\bm{w}})+\psi(\hat{\bm{w}})$ over parameter vectors ${\bm{w}},$ where $l_{i}$ is a univariate Lipschitz-continuous function and $\psi$ is proximable, addressing $W_{1}$-DRO examples from Section 3.3. Due to the discussion from Section 3.4, effects of data contamination can be reduced to an inexact oracle ${\mathcal{O}}({\bm{\beta}},\{{\bm{x}}_{i}\})$, as stated in Assumption 1.4.
For conciseness of notation, let $l_{i}\equiv\ell_{y_{i}}$ and let $g_{i}\equiv l_{i}^{*}$ be its convex conjugate. The pseudocode for our algorithm is provided in Algorithm 1. Algorithm 1 can be seen as a variant of primal-dual hybrid gradient (PDHG) [chambolle2011first], but with two key differences addressed by our analysis: (1) the “extrapolated gradient vector” used in the computation of ${\bm{w}}_{k}$ (see Line 6 in Algorithm 1) is accessed using a bounded-error oracle ${\mathcal{O}}({\bm{\beta}},\{{\bm{x}}_{i}\})$, and (2) the amount of primal regularization (the quadratic term $\|{\bm{u}}-{\bm{w}}_{k-1}\|_{2}^{2}$ in the definition of ${\bm{w}}_{k}$) is scaled by $c_{k}$, which monotonically decreases over iterations $k$. The item (2) is crucial for establishing an error guarantee that scales with $\|{\bm{w}}_{0}-{\bm{w}}_{*}\|_{2}$ in place of the maximum distance of iterates to ${\bm{w}}_{*},$ which may be (much) larger. We also highlight that the inexact oracle is used only in the computation of primal updates, whereas the dual updates (in Line 8) are computed using corrupted samples.
1 Input $\epsilon$, corrupted samples $\{({\bm{x}}_{i}^{c},y_{i}^{c})\}_{i=1}^{N}$, $\sigma$, ${\bm{w}}_{0}={\bm{0}},\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$, estimation guarantee $\Delta$
@@ 4       $A_{k}=a_{k}+A_{k-1},c_{k}=c_{k-1}-1/T$
1 Input & Initialization stable samples $\{({\bm{x}}_{i}^{s},y_{i}^{s})\}_{i=1}^{N}$, same other parameters as in Algorithm 1
@@
Let ${\bm{w}}^{*}\in\operatorname*{arg\,min}_{{\bm{w}}\in\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}\ell_{y_{i}}({\bm{x}}^{s}_{i}\cdot{\bm{w}})+\psi({\bm{w}})$, where $\{{\bm{x}}^{s}_{i}\}_{i=1}^{N}$ are the stable covariates. Under Assumptions 1.3 and 1.4, Algorithm 2 outputs $\hat{\bm{w}}_{T}$ that satisfies

### v2

@@ 3.3 Equivalent DRO Reformulation
In this work, we focus on convex Lipschitz-continuous functions that can be written as a composition of a high-dimensional affine function and a univariate nonlinear function. In particular, if for labeled data $({\bm{x}},y)$ the loss function $\ell({\bm{w}};{\bm{x}},y)$ takes the form of either $\phi_{1}({\bm{w}}\cdot{\bm{x}}-y)$ [^9] More generally, for loss functions of the form $\phi_{1}({\bm{w}}\cdot{\bm{x}}-y)$, we can also allow for distributional shifts in the label $y$: when $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|[{\bm{x}}^{\top},y]^{\top}-[{\bm{x}}^{\prime\top},y^{\prime}]^{\top}\|_{r}$, we replace the regularization $\|{\bm{w}}\|_{s}$ by $\|[{\bm{w}}^{\top},-1]^{\top}\|_{s}$. For conciseness, in this paper we only consider distributional shifts in the covariates $\bm{x}$}. or $\phi_{2}(y{\bm{w}}\cdot{\bm{x}})$ for univariate functions $\phi_{1}$ and $\phi_{2}$ of Lipschitz modulus $\zeta$ and the cost function is $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}+\chi(y-y^{\prime})$ for some $r\in[1,+\infty]$, where $\chi(0)=0$ and is $\infty$ elsewhere, then [SKE19a, Theorem 4 and Remark 18 ] implies that (1) is equivalent to
@@
where $1/r+1/s=1.$ Thus, our problem reduces to solving (6) with sample access to data-contaminated distribution ${\mathbb{P}}_{\text{train}}$} (cf. (2)). Due to standard uniform convergence results, it suffices to robustly solve an empirical version of (6) as in problem (P).
We give a few examples of reformulating Equation˜1 as Equation˜6. Recall that the cost function is $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}+\chi(y-y^{\prime})$; in the following, we drop the superscript $c$, since the context is clear. The convex conjugate of these loss functions and their efficient proximal operator evaluation per Definition˜2.1 can be found in [Bec17a, Sections 4.4.16 and 6.9 ].
@@
Our results directly apply to both Example˜3.5 and Example˜3.6, thanks to the finite-sum structure of their empirical loss functions. Below, we also include the example of $W_{2}$-robust linear regression, which does not exhibit the same finite-sum loss structure (notice the square-root in (7)), but to which our results apply after a simple transformation; see Remark˜4.6.
@@ 3.4 Inexact Hybrid Gradient Oracle
Finally, we argue that under standard distributional assumptions (stated as Assumption˜3.8), we can assume access to a bounded-error oracle for data drawn from ${\mathbb{P}}_{0}.$
@@
[…]
We prove the following key technical lemma that enables the robust estimation of inexact hybrid gradients for all iterations, which is a nontrivial task because iterates depend on previous iterations.
@@
Proof.
Let ${\bm{z}}_{i}=\beta_{i}{\bm{x}}_{i}$ for each $i\in[N]$. Define the empirical mean of the scaled set as ${\bm{\mu}}_{\bm{z}}=\frac{1}{N}\sum_{i=1}^{N}{\bm{z}}_{i}$, and its second moment matrix as ${\bm{M}}_{\bm{z}}=\frac{1}{N}\sum_{i=1}^{N}{\bm{z}}_{i}{\bm{z}}_{i}^{\top}$. The empirical covariance matrix of the scaled set is ${\bm{\mathrm{Cov}}}(\{{\bm{z}}_{i}\})={\bm{M}}_{\bm{z}}-{\bm{\mu}}_{\bm{z}}{\bm{\mu}}_{\bm{z}}^{\top}$. Since ${\bm{\mu}}_{\bm{z}}{\bm{\mu}}_{\bm{z}}^{\top}$ is Positive Semi-Definite (PSD), we have ${\bm{\mathrm{Cov}}}(\{{\bm{z}}_{i}\})\preceq{\bm{M}}_{\bm{z}}$, which implies $\|{\bm{\mathrm{Cov}}}(\{{\bm{z}}_{i}\})\|_{\operatorname{op}}\leq\|{\bm{M}}_{\bm{z}}\|_{\operatorname{op}}$.
Now, we relate ${\bm{M}}_{\bm{x}}$ to the covariance matrix ${\bm{\mathrm{Cov}}}(\{{\bm{x}}_{i}\})$ of the original set: ${\bm{M}}_{\bm{x}}={\bm{\mathrm{Cov}}}(\{{\bm{x}}_{i}\})+{\bm{\mu}}{\bm{\mu}}^{\top}$. Thus, $\|{\bm{M}}_{\bm{x}}\|_{\operatorname{op}}\leq\|{\bm{\mathrm{Cov}}}(\{{\bm{x}}_{i}\})\|_{\operatorname{op}}+\|{\bm{\mu}}{\bm{\mu}}^{\top}\|_{\operatorname{op}}$. Since ${\bm{\mu}}{\bm{\mu}}^{\top}$ is a rank-one PSD matrix, $\|{\bm{\mu}}{\bm{\mu}}^{\top}\|_{\operatorname{op}}=\|{\bm{\mu}}\|_{2}^{2}$. By Fact˜2.6, we know $\|{\bm{\mathrm{Cov}}}(\{{\bm{x}}_{i}\})\|_{\operatorname{op}}\leq\sigma^{2}$, so $\|{\bm{M}}_{\bm{x}}\|_{\operatorname{op}}\leq\sigma^{2}+\|{\bm{\mu}}\|_{2}^{2}$.
Combining the inequalities, we conclude that $\|{\bm{\mathrm{Cov}}}(\{{\bm{z}}_{i}\})\|_{\operatorname{op}}\leq\|{\bm{M}}_{\bm{z}}\|_{\operatorname{op}}\leq\zeta^{2}(\sigma^{2}+\|{\bm{\mu}}\|_{2}^{2})$. Another application of Fact˜2.6 completes the proof. ∎
This lemma implies that we can use Lemma˜3.9 to construct an oracle that satisfies Assumption˜1.4 via robust mean estimation algorithms described in Section 2.
@@
On input an arbitrary $3\zeta$-uniformly bounded sequence $(\beta^{i})_{i=1}^{N}$ and a $2\epsilon$-corrupted version of an $(O(\epsilon),O(\sqrt{\epsilon}))$-stable set $\{{\bm{x}}_{i}\}_{i=1}^{N}$ with respect to variance $\sigma^{2}$, $\mathsf{RobustMeanEstimation}$ (Algorithm˜3) outputs $\hat{{\bm{z}}}$ such that $\|\hat{\bm{z}}-\frac{1}{N}\sum_{i=1}^{N}\beta^{i}{\bm{x}}_{i}\|_{2}=O(\sigma\zeta\sqrt{\epsilon})$.
We remark on the parameter scaling in the error bound of Proposition˜3.10. Recall that the term $O(\sigma\sqrt{\epsilon})$ is the information-theoretically optimal error for robustly estimating the mean of a distribution with covariance bounded by $\sigma^{2}I$ (even in one dimension); see Fact˜4.7 below for a precise statement. Our result shows that this error guarantee can be extended to the more complex setting of estimating a weighted mean of covariates; the additional factor of $\zeta$ appears as a result of the scaling. This oracle error of $\Delta=O(\zeta\sigma\sqrt{\epsilon})$ propagates to our final optimality gap, which Lemma˜4.8 proves to be information-theoretically optimal.
@@ 4 Algorithm and Analysis
[…]
For conciseness of notation, let $l_{i}\equiv\ell_{y_{i}}$ and let $g_{i}\equiv l_{i}^{*}$ be its convex conjugate. The pseudocode for our algorithm is provided in Algorithm˜1. Algorithm˜1 can be seen as a variant of primal-dual hybrid gradient (PDHG) [CP11a], but with two key differences addressed by our analysis: (1) the “extrapolated gradient vector” used in the computation of ${\bm{w}}_{k}$ (see Line 6 in Algorithm˜1) is accessed using a bounded-error oracle ${\mathcal{O}}({\bm{\beta}},\{{\bm{x}}_{i}\})$, and (2) the amount of primal regularization (the quadratic term $\|{\bm{u}}-{\bm{w}}_{k-1}\|_{2}^{2}$ in the definition of ${\bm{w}}_{k}$) is scaled by $c_{k}$, which monotonically decreases over iterations $k$. The item (2) is crucial for establishing an error guarantee that scales with $\|{\bm{w}}_{0}-{\bm{w}}_{*}\|_{2}$ in place of the maximum distance of iterates to $\bm{w}$$_{*}$, which may be (much) larger. We also highlight that the inexact oracle is used only in the computation of primal updates, whereas the dual updates (in Line 8) are computed using corrupted samples.
1Input $\epsilon$, corrupted samples $\{({\bm{x}}_{i}^{c},y_{i}^{c})\}_{i=1}^{N}$, $\sigma$, ${\bm{w}}_{0}={\bm{0}},\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$, estimation guarantee $\Delta$
@@ 4 $A_{k}=a_{k}+A_{k-1},c_{k}=c_{k-1}-1/T$, where $a_{k}=\sqrt{N}/\sigma$
1Input & Initialization stable samples $\{({\bm{x}}_{i}^{s},y_{i}^{s})\}_{i=1}^{N}$, same other parameters as in Algorithm˜1
@@
Let ${\bm{w}}^{*}\in\operatorname*{arg\,min}_{{\bm{w}}\in\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}\ell_{y_{i}}({\bm{x}}^{s}_{i}\cdot{\bm{w}})+\psi({\bm{w}})$, where $\{{\bm{x}}^{s}_{i}\}_{i=1}^{N}$ are the stable covariates. Under Assumptions˜1.3 and 1.4, Algorithm˜2 outputs $\hat{\bm{w}}_{T}$ that satisfies